# Characterizations of voting rules based on majority margins

### v2

@@ 4.2 Neutral Indifference
A head-to-head voting rule on the domain of all profiles (resp. on LOBI) is C2 if and only if it satisfies Neutral Indifference.
@@
The left-to-right direction is immediate. From right to left, assume a head-to-head voting rule $F$ satisfies Neutral Indifference, and consider profiles $\mathbf{P},\mathbf{P}^{\prime}$ such that #$\#_{\mathbf{P}}=\#_{\mathbf{P}^{\prime}}$. If in addition $|V(\mathbf{P})|=|V(\mathbf{P}^{\prime})|$, then $\mathcal{H}(\mathbf{P})=\mathcal{H}(\mathbf{P}^{\prime})$, so by the head-to-head assumption, $F(\mathbf{P})=F(\mathbf{P}^{\prime})$. If one profile, say $\mathbf{P}$, has fewer voters, then add voters with fully indifferent ballots until we obtain a profile $\mathbf{P}^{+}$ with the same number of voters as $\mathbf{P}^{\prime}$. Then by Neutral Indifference, $F(\mathbf{P})=F(\mathbf{P}^{+})$, and by the head-to-head assumption, $F(\mathbf{P}^{+})=F(\mathbf{P}^{\prime})$, so $F(\mathbf{P})=F(\mathbf{P}^{\prime})$, which shows that $F$ is C2.∎
@@
Note that the Pareto rule violates Neutral Indifference, while the version of Minimax based on winning votes satisfies it. It is also easy to see that Neutral Indifference and Nonlinear Neutral Reversal are independent, as follows.
Fact 4.5.
@@ 4.3 Proof of characterization
The left-to-right direction of the ‘if and only if’ is obvious. The right-to-left direction is an immediate corollary of the following lemma stating that whenever two profiles have the same margin information, they can be transformed in the ways described by Nonlinear Neutral Reversal and Neutral Indifference to profiles having the same $\mathcal{H}$ information as defined in (3), in which case a head-to-head rule must yield the same output for the profiles.
Lemma 4.7.
For any profiles $\mathbf{P},\mathbf{P}^{\prime}$, if $\mathcal{M}_{\mathbf{P}}=\mathcal{M}_{\mathbf{P}^{\prime}}$, then there are profiles $\mathbf{Q}$ and $\mathbf{Q}^{\prime}$ such that:
@@
    $\mathbf{Q}$ (resp. $\mathbf{Q}^{\prime}$) is obtained by adding (possibly zero) (possibly nonlinear) reversal pairs of voters or voters with the empty ballot to $\mathbf{P}$ (resp. $\mathbf{P}^{\prime}$);
@@
    $\mathcal{H}(\mathbf{Q})=\mathcal{H}(\mathbf{Q^{\prime}})$.
Proof.
Let $\mathbf{P}$ and $\mathbf{P}^{\prime}$ satisfying the requirement be given with $X:=X(\mathbf{P})=X(\mathbf{P}^{\prime})$. Recall that $X$ is finite. For any doubleton $D\subseteq X$, we define two profiles $\mathbf{A}_{D}$ and $\mathbf{A}^{\prime}_{D}$ as follows. Let $a$ and $b$ be the smaller and the greater elements of $D$, respectively, according to the alphabetic ordering on $X$. Then, let
@@
Note that since by assumption $\mathcal{M}_{\mathbf{P}}=\mathcal{M}_{\mathbf{P}^{\prime}}$, $d_{D}$ is also equal to #$\#_{\mathbf{P}^{\prime}}(b,a)-\#_{\mathbf{P}}(b,a)$. So, as long as $D=\{x,y\}$, $d_{D}=\#_{\mathbf{P}^{\prime}}(x,y)-\#_{\mathbf{P}}(x,y)$.
@@
  • If $d_{D}<0$, switch the role of $\mathbf{A}_{D}$ and $\mathbf{A}^{\prime}_{D}$ in the above construction while using $-d_{D}$ pairs of fresh voters for each profile. That is, $\mathbf{A}_{D}$ will use $-d_{D}$ pairs of $T$ and $T^{-1}$ ballots while $\mathbf{A}^{\prime}_{D}$ will use $-d_{D}$ pairs of $L$ and $L^{-1}$.
  • It is instructive to do a little counting here.
    • For any $x,y\in X\setminus D$, plainly #$\#_{\mathbf{A}_{D}}(x,y)=\#_{\mathbf{A}^{\prime}_{D}}(x,y)=0$ since they are always in the same equivalence class.
    • For any $x\in D$ and $y\in X\setminus D$, we have #$\#_{\mathbf{A}_{D}}(x,y)=\#_{\mathbf{A}^{\prime}_{D}}(x,y)=d_{D}$ (by $L$ and $T$) and also #$\#_{\mathbf{A}_{D}}(y,x)=\#_{\mathbf{A}^{\prime}_{D}}(y,x)=d_{D}$ (by $L^{-1}$ and $T^{-1}$).
@@
      #$\#_{\mathbf{A}_{D}}(a,b)=\#_{\mathbf{A}_{D}}(b,a)=d_{D}\mbox{ while }\#_{\mathbf{A}^{\prime}_{D}}(a,b)=\#_{\mathbf{A}^{\prime}_{D}}(b,a)=0.$
@@
      #$\#_{\mathbf{A}_{D}}(a,b)=\#_{\mathbf{A}_{D}}(b,a)=0\mbox{ while }\#_{\mathbf{A}^{\prime}_{D}}(a,b)=\#_{\mathbf{A}^{\prime}_{D}}(b,a)=-d_{D}.$
@@
      #$\#_{\mathbf{A}_{D}}(a,b)=\#_{\mathbf{A}_{D}}(b,a)=0\mbox{ and }\#_{\mathbf{A}^{\prime}_{D}}(a,b)=\#_{\mathbf{A}^{\prime}_{D}}(b,a)=0.$
@@
      #$\#_{\mathbf{A}_{D}}(a,b)-\#_{\mathbf{A}^{\prime}_{D}}(a,b)=\#_{\mathbf{A}_{D}}(b,a)-\#_{\mathbf{A}^{\prime}_{D}}(b,a)=d_{D}.$
@@
    #$\#_{\mathbf{A}_{D}}(x,y)-\#_{\mathbf{A}^{\prime}_{D}}(x,y)=\begin{cases}0&\text{ if }\{x,y\}\not=D\\ d_{D}&\text{ if }\{x,y\}=D.\end{cases}$
  • Also notice that both $\mathbf{A}_{D}$ and $\mathbf{A}^{\prime}_{D}$ consist solely of possibly non-linear reversal pairs.
Now let $\mathbf{Q}$ (resp. $\mathbf{Q}^{\prime}$) be the disjoint union of $\mathbf{P}$ (resp. $\mathbf{P}^{\prime}$) and $\mathbf{A}_{D}$ (resp. $\mathbf{A}^{\prime}_{D}$) for all doubletons $D\subseteq X$. Then #$\#_{\mathbf{Q}}=\#_{\mathbf{Q}^{\prime}}$. This is because, for any $x\not=y\in X$,
#$\#_{\mathbf{Q}}(x,y)-\#_{\mathbf{Q}^{\prime}}(x,y)$
=(#$=\left(\#_{\mathbf{P}}(x,y)+\sum_{D\subseteq X,|D|=2}\#_{\mathbf{A}_{D}}(x,y)\right)-\left(\#_{\mathbf{P}^{\prime}}(x,y)+\sum_{D\subseteq X,|D|=2}\#_{\mathbf{A}^{\prime}_{D}}(x,y)\right)$
=#$=\#_{\mathbf{P}}(x,y)-\#_{\mathbf{P}^{\prime}}(x,y)+\sum_{D\subseteq X,|D|=2}(\#_{\mathbf{A}_{D}}(x,y)-\#_{\mathbf{A}^{\prime}_{D}}(x,y))$
@@
$=0.$
At this point, it may be that $|V(\mathbf{Q})|\not=|V(\mathbf{Q}^{\prime})|$. However, we can simply add to the smaller profile enough voters with the empty ballot so that $\mathbf{Q}$ and $\mathbf{Q}^{\prime}$ have the same size. This step clearly does not change #$\#_{\mathbf{Q}}$ and #$\#_{\mathbf{Q}^{\prime}}$, so they are still equal. ∎
@@
Combining Theorem 4.6 and Proposition 4.4, we obtain the implications shown in Figure 4.
@@ 5 Conclusion
It is an interesting open question whether other invariance properties can be proved equivalent to combinations of natural, normatively plausible axioms. For example, can the class of head-to-head voting rules be characterized in terms of Preferential Equality together with some independent axioms? While the class of head-to-head rules is larger than that of margin-based rules, we can also go in the other direction and consider more restricted classes. For example, many of the well-known margin-based rules (recall Footnote 1) satisfy the stronger property of ordinal margin invariance (see Holliday 2024), according to which we do not need to know the exact sizes of margins but only whether the margin of $x$ over $y$ is greater than that of $x^{\prime}$ over $y^{\prime}$. Or one could go all the way to Fishburn’s (1977) class of C1 rules, according to which all we need to know about the margins is their signs (positive, zero, or negative). It is doubtful that increasingly strong invariance properties will continue to have normatively compelling equivalent axiomatizations, since at some point (arguably by the time we reach C1) the invariance properties appear clearly too strong. Indeed, the project of seeking normatively perspicuous re-axiomatizations of invariance properties can guide us toward the boundaries between desirable and undesirable invariance properties for voting rules.
@@ Appendix A Appendix: Allowing Noncomparability
We also change our definition of a profile to mean a function $\mathbf{R}:V\to\wp(X^{2})$ where $V$ is a nonempty finite subset of $\mathcal{V}$, also denoted by $V(\mathbf{R})$, and $X$ is a nonempty finite subset of $\mathcal{X}$, also denoted by $X(\mathbf{R})$. For any profile $\mathbf{R}$, its head-to-head counting functions #$\#P_{\mathbf{R}},\#I_{\mathbf{R}},\#N_{\mathbf{R}}:X(\mathbf{R})^{2}\to\mathbb{N}$ are defined by
  • #$\#P_{\mathbf{R}}(x,y)=|\{i\in V(\mathbf{R})\mid xP(\mathbf{R}(i))y\}|$,
  • #$\#I_{\mathbf{R}}(x,y)=|\{i\in V(\mathbf{R})\mid xI(\mathbf{R}(i))y\}|$, and
  • #$\#N_{\mathbf{R}}(x,y)=|\{i\in V(\mathbf{R})\mid xN(\mathbf{R}(i))y\}|$.
Then the total head-to-head information $\mathcal{H}(\mathbf{R})$ of $\mathbf{R}$ is simply the tuple (#$(\#P_{\mathbf{R}},\#I_{\mathbf{R}},\#N_{\mathbf{R}})$ collecting the three counting functions. The margin function is defined as before:
@@
    If the domain of $F$ is $\mathsf{WOSN}$, then $F$ is margin-based if and only if $F$ satisfies Tiebreaking Compensation, Comparable Compensation, and Neutral Reversal.
@@
If a profile contains a ranking with nonempty side-noncomparability class $S$, duplicate the profile, so we have two copies of the ranking, and then linearize $S$ in opposite ways above and below the other candidates. By Homogeneity and Comparable Compensation, this does not change the output of $F$. In this way, we can change any two initial profiles into profiles with no side-noncomparability. Then for part 1 we reason exactly as in the proof of Theorem 2.9 using Preferential Equality, and for part 2 we reason exactly as in the proof of Theorem 3.8.2 using Tiebreaking Compensation.∎

### v3

@@ 3.2 Neutral Indifference
A head-to-head voting rule on the domain of all profiles (resp. on $\mathsf{LOBI}$) is C2 if and only if it satisfies Neutral Indifference.
@@
The left-to-right direction is immediate. From right to left, assume a head-to-head voting rule $F$ satisfies Neutral Indifference, and consider profiles $\mathbf{P},\mathbf{P}^{\prime}$ such that $_{\mathbf{P}}=\#_{\mathbf{P}^{\prime}}$. If in addition $|V(\mathbf{P})|=|V(\mathbf{P}^{\prime})|$, then $\mathcal{H}(\mathbf{P})=\mathcal{H}(\mathbf{P}^{\prime})$, so by the head-to-head assumption, $F(\mathbf{P})=F(\mathbf{P}^{\prime})$. If one profile, say $\mathbf{P}$, has fewer voters, then add voters with fully indifferent ballots until we obtain a profile $\mathbf{P}^{+}$ with the same number of voters as $\mathbf{P}^{\prime}$. Then by Neutral Indifference, $F(\mathbf{P})=F(\mathbf{P}^{+})$, and by the head-to-head assumption, $F(\mathbf{P}^{+})=F(\mathbf{P}^{\prime})$, so $F(\mathbf{P})=F(\mathbf{P}^{\prime})$, which shows that $F$ is C2.∎
@@
Note that the strict Pareto rule violates Neutral Indifference, while the version of Minimax based on winning votes satisfies it.
3.3 Proof of characterization
Theorem 3.8.
Let $F$ be a voting rule satisfying Homogeneity. [^12] In fact, for this theorem, it suffices to assume the weaker form of Homogeneity that only requires $F(\mathbf{P})=F(2\mathbf{P})$.
  1. 1.
    If the domain of $F$ is $\mathsf{LOBI}$, then $F$ is margin-based if and only if $F$ satisfies Preferential Equality, Tiebreaking Compensation, and Neutral Indifference.
  2. 2.
    If the domain of $F$ is the domain of all profiles, then $F$ is margin-based if and only if $F$ satisfies Tiebreaking Compensation and Neutral Indifference.
Proof.
If a profile contains a ranking with a tie, duplicate the profile, so we have two copies of the ranking, and then linearize the previously tied part in opposite ways. By Homogeneity and Tiebreaking Compensation, this does not change the output of $F$. In this way, we can change any two initial profiles into linear profiles. Then we reason exactly as in the proof of Theorem 2.9 using Preferential Equality, which is available in part 2 thanks to Lemma 3.3, and Neutral Reversal, which is a consequence of Neutral Indifference (applied twice) together with Tiebreaking Compensation.∎
Remark 3.9.
In part 1 of Theorem 3.8, we can replace Tiebreaking Compensation by what could be called Pure Tiebreaking Compensation: if two voters have the same ranking containing a tie $Y$, then changing the profile so that each voter linearizes $Y$ in opposite ways does not change the output of the voting rule. For in the proof of Theorem 3.8.1, we apply Tiebreaking Compensation to two copies of the same ranking. We could also replace Tiebreaking Compensation with Pure Tiebreaking Compensation in part 2 of Theorem 3.8, provided we add Preferential Equality, since unlike Tiebreaking Compensation, Pure Tiebreaking Compensation does not imply Preferential Equality for voting rules on the domain of all profiles. [^13] Given $Y\subsetneq X$ and rankings $P$ and $P^{\prime}$ of $X$, say that $P$ and $P^{\prime}$ agree outside $Y$, denoted $P\sim_{Y}P^{\prime}$, if (a) for all $x\in X\setminus Y$, either $xPy$ for all $y\in Y$ or $yPx$ for all $y\in Y$, and either $xP^{\prime}y$ for all $y\in Y$ or $yP^{\prime}x$ for all $y\in Y$, and (b) $P\setminus Y^{2}=P^{\prime}\setminus Y^{2}$. Let $\mathbb{P}$ be the set of profiles $\mathbf{P}$ such that there are two voters whose rankings $P$ and $P^{\prime}$ agree outside some $Y\subsetneq X(\mathbf{P})$ or are such that $P^{\prime}=P^{-1}$. Let $F$ be the voting rule such that if $\mathbf{P}\in\mathbb{P}$, then $F(\mathbf{P})=X(\mathbf{P})$, and otherwise $F(\mathbf{P})$ is the set of Plurality winners. Observe that if $\mathbf{Q}^{\prime}$ is obtained from $\mathbf{Q}$ by an application of Pure Tiebreaking Compensation, then $\mathbf{Q},\mathbf{Q}^{\prime}\in\mathbb{P}$, so $F$ satisfies Pure Tiebreaking Compensation. However, $F$ violates Preferential Equality for the profile $\mathbf{P}$ with just two ballots, abcdeabcde and dabecdabec, since switching $ab$ to $ba$ in the first ballot vs. the second ballot leads to profiles that are not in $\mathbb{P}$ and have different Plurality winners.
Fact 3.10.
4 Characterization modulo head-to-head
We now return to the distinction introduced in Section 1 between margin-based voting rules and the larger class of head-to-head voting rules. We already know as a consequence of Theorem 3.8.2 that a homogeneous, head-to-head voting rule on the domain of all profiles is margin-based if and only if it satisfies Tiebreaking Compensation and Neutral Indifference. However, once we assume that the voting rule is head-to-head, we can provide an alternative axiomatization of margin-based rules using only Neutral Indifference and a strengthening of Neutral Reversal in the context of possibly nonlinear ballots called Nonlinear Neutral Reversal, introduced in the next subsection.
In this section, we work with voting rules on the domain of all profiles (for $\mathsf{LOBI}$, we refer the reader back to Theorem 3.8.1). For notational convenience, where $S_{1},\dots,S_{n}$ partition $X$,
$S_{1}>S_{2}>\dots>S_{n}$
4.1 Nonlinear Neutral Reversal
The new axiom we will use in our characterization of margin-based rules relative to the class of head-to-head rules is a strengthening of Neutral Reversal. It applies not only when two voters submit reversed linear orders but also when two voters submit reversed rankings with ties.
Definition 4.1.
Fact 4.2.
Let $F$ be the version of the Borda Count for profiles of strict weak orders defined as follows: from a strict weak order $P$, a candidate $x$ receives $|\{y\in X(\mathbf{P})\mid(x,y)\in P\}|$ points, and the candidates with the most points win. Then adding the pair of rankings $\{a\}>X(\mathbf{P})\setminus\{a\}$ and $X(\mathbf{P})\setminus\{a\}>\{a\}$ to a profile gives $|X(\mathbf{P})|-1$ points to $a$ and $1$ point to everyone else, which can change the set of winners, so $F$ does not satisfy Nonlinear Neutral Reversal. Yet it is easy to see that $F$ does satisfy Neutral Reversal.
It is also easy to see that Neutral Indifference and Nonlinear Neutral Reversal are independent, as follows.
Fact 4.3.
@@ 4.2 Proof of characterization
The left-to-right direction of the ‘if and only if’ is obvious. The right-to-left direction is an immediate corollary of Proposition 3.7 characterizing C2 voting rules among head-to-head voting rules using Neutral Indifference and the following proposition characterizing margin-based voting rules among C2 voting rules using Nonlinear Neutral Reversal.
Proposition 4.5.
A C2 voting rule on the domain of all profiles is margin-based if and only if it satisfies Nonlinear Neutral Reversal.
Proof.
For the non-trivial direction, it is enough to show the following: for any profiles $\mathbf{P},\mathbf{P}^{\prime}$, if $\mathcal{M}_{\mathbf{P}}=\mathcal{M}_{\mathbf{P}^{\prime}}$, then there are profiles $\mathbf{Q}$ and $\mathbf{Q}^{\prime}$ such that:
@@
    $\mathbf{Q}$ (resp. $\mathbf{Q}^{\prime}$) is obtained by adding (possibly zero) (possibly nonlinear) reversal pairs of voters to $\mathbf{P}$ (resp. $\mathbf{P}^{\prime}$);
@@
    $_{\mathbf{Q}}=\#_{\mathbf{Q}^{\prime}}$.
For then by part 1 and Nonlinear Neutral Reversal, $F(\mathbf{P})=F(\mathbf{Q})$ and $F(\mathbf{P}^{\prime})=F(\mathbf{Q}^{\prime})$, and by part 2 and C2, $F(\mathbf{Q})=F(\mathbf{Q}^{\prime})$, so $F(\mathbf{P})=F(\mathbf{P}^{\prime})$, showing that $F$ is margin-based.
Let $\mathbf{P}$ and $\mathbf{P}^{\prime}$ satisfying the requirement be given with $X:=X(\mathbf{P})=X(\mathbf{P}^{\prime})$. Recall that $X$ is finite. For any doubleton $D\subseteq X$, we define two profiles $\mathbf{A}_{D}$ and $\mathbf{A}^{\prime}_{D}$ as follows. Let $a$ and $b$ be the smaller and the greater elements of $D$, respectively, according to the alphabetic ordering of $X$. Then let
@@
Note that since by assumption $\mathcal{M}_{\mathbf{P}}=\mathcal{M}_{\mathbf{P}^{\prime}}$, $d_{D}$ is also equal to $_{\mathbf{P}^{\prime}}(b,a)-\#_{\mathbf{P}}(b,a)$. So, as long as $D=\{x,y\}$, $d_{D}=\#_{\mathbf{P}^{\prime}}(x,y)-\#_{\mathbf{P}}(x,y)$.
@@
  • If $d_{D}<0$, switch the role of $\mathbf{A}_{D}$ and $\mathbf{A}^{\prime}_{D}$ in the above construction while using $d_{D}$ pairs of fresh voters for each profile. That is, $\mathbf{A}_{D}$ will use $d_{D}$ pairs of $T$ and $T^{-1}$ ballots while $\mathbf{A}^{\prime}_{D}$ will use $d_{D}$ pairs of $L$ and $L^{-1}$.
  • It is instructive to do a little counting here:
    • For any $x,y\in X\setminus D$, plainly $_{\mathbf{A}_{D}}(x,y)=\#_{\mathbf{A}^{\prime}_{D}}(x,y)=0$ since they are always in the same equivalence class.
    • For any $x\in D$ and $y\in X\setminus D$, we have $_{\mathbf{A}_{D}}(x,y)=\#_{\mathbf{A}^{\prime}_{D}}(x,y)=|d_{D}|$ (by $L$ and $T$) and also $_{\mathbf{A}_{D}}(y,x)=\#_{\mathbf{A}^{\prime}_{D}}(y,x)=|d_{D}|$ (by $L^{-1}$ and $T^{-1}$).
@@
      $_{\mathbf{A}_{D}}(a,b)=\#_{\mathbf{A}_{D}}(b,a)=d_{D}\mbox{ while }\#_{\mathbf{A}^{\prime}_{D}}(a,b)=\#_{\mathbf{A}^{\prime}_{D}}(b,a)=0.$
@@
      $_{\mathbf{A}_{D}}(a,b)=\#_{\mathbf{A}_{D}}(b,a)=0\mbox{ while }\#_{\mathbf{A}^{\prime}_{D}}(a,b)=\#_{\mathbf{A}^{\prime}_{D}}(b,a)=-d_{D}.$
@@
[…]
@@
      $_{\mathbf{A}_{D}}(a,b)-\#_{\mathbf{A}^{\prime}_{D}}(a,b)=\#_{\mathbf{A}_{D}}(b,a)-\#_{\mathbf{A}^{\prime}_{D}}(b,a)=d_{D}.$
@@
    $\#_{\mathbf{A}_{D}}(x,y)-\#_{\mathbf{A}^{\prime}_{D}}(x,y)=\begin{cases}0&\text{ if }\{x,y\}\not=D\\ d_{D}&\text{ if }\{x,y\}=D.\end{cases}$
  • Also notice that both $\mathbf{A}_{D}$ and $\mathbf{A}^{\prime}_{D}$ consist solely of possibly nonlinear reversal pairs.
Now let $\mathbf{Q}$ (resp. $\mathbf{Q}^{\prime}$) be the disjoint union of $\mathbf{P}$ (resp. $\mathbf{P}^{\prime}$) and $\mathbf{A}_{D}$ (resp. $\mathbf{A}^{\prime}_{D}$) for all doubletons $D\subseteq X$. Then $_{\mathbf{Q}}=\#_{\mathbf{Q}^{\prime}}$. This is because, for any $x\not=y\in X$,
$\#_{\mathbf{Q}}(x,y)-\#_{\mathbf{Q}^{\prime}}(x,y)$
$=\left(\#_{\mathbf{P}}(x,y)+\sum_{D\subseteq X,|D|=2}\#_{\mathbf{A}_{D}}(x,y)\right)-\left(\#_{\mathbf{P}^{\prime}}(x,y)+\sum_{D\subseteq X,|D|=2}\#_{\mathbf{A}^{\prime}_{D}}(x,y)\right)$
$=\#_{\mathbf{P}}(x,y)-\#_{\mathbf{P}^{\prime}}(x,y)+\sum_{D\subseteq X,|D|=2}(\#_{\mathbf{A}_{D}}(x,y)-\#_{\mathbf{A}^{\prime}_{D}}(x,y))$
@@
$=0.\qed$
@@
Combining Proposition 3.7 and Proposition 4.5, we obtain the implications shown in Figure 4.
@@ 5 Conclusion
It is an interesting open question whether other invariance properties can be proved equivalent to combinations of natural, normatively significant axioms. For example, can the class of head-to-head voting rules be characterized in terms of Preferential Equality together with some independent axioms? [^15] To apply the strategy used in this paper, we would need to show that if two profiles share the same head-to-head information, then there is a way to turn them into the same profile using operations preserving the head-to-head information, and in particular, the number of voters. This seems to require a detailed analysis of when and how the same margin matrix can be realized by multiple profiles of the same size. While the class of head-to-head rules is larger than that of margin-based rules, we can also go in the other direction and consider more restricted classes. For example, many of the well-known margin-based rules (recall Footnote 1) satisfy the stronger property of ordinal margin invariance (see Holliday 2024), according to which we do not need to know the exact sizes of margins but only whether the margin of $x$ over $y$ is greater than that of $x^{\prime}$ over $y^{\prime}$. Or one could go all the way to Fishburn’s (1977) class of C1 rules, according to which all we need to know about the margins is their signs (positive, zero, or negative). It is doubtful that increasingly strong invariance properties will continue to have normatively compelling equivalent axiomatizations, since at some point (arguably by the time we reach C1) the invariance properties appear clearly too strong. Indeed, the project of seeking normatively perspicuous re-axiomatizations of invariance properties may guide us toward the boundaries between desirable and undesirable invariance properties for voting rules.
Acknowledgements
We thank the audiences at the COMSOC Video Seminar in January 2025 and at the Special Session on the Mathematics of Decisions, Elections, and Games at the Joint Mathematics Meeting 2025, as well as the anonymous referees for Social Choice and Welfare for their helpful feedback.
@@ Appendix A Appendix: Allowing Noncomparability
We also change our definition of a profile to mean a function $\mathbf{R}:V\to\wp(X^{2})$ where $V$ is a nonempty finite subset of $\mathcal{V}$, also denoted by $V(\mathbf{R})$, and $X$ is a nonempty finite subset of $\mathcal{X}$ with $|X|\geq 2$, also denoted by $X(\mathbf{R})$. For any profile $\mathbf{R}$, its head-to-head counting functions $P_{\mathbf{R}},\#I_{\mathbf{R}},\#N_{\mathbf{R}}:X(\mathbf{R})^{2}\to\mathbb{N}$ are defined by
  • $P_{\mathbf{R}}(x,y)=|\{i\in V(\mathbf{R})\mid xP(\mathbf{R}(i))y\}|$,
  • $I_{\mathbf{R}}(x,y)=|\{i\in V(\mathbf{R})\mid xI(\mathbf{R}(i))y\}|$, and
  • $N_{\mathbf{R}}(x,y)=|\{i\in V(\mathbf{R})\mid xN(\mathbf{R}(i))y\}|$.
Then the total head-to-head information $\mathcal{H}(\mathbf{R})$ of $\mathbf{R}$ is simply the tuple $(\#P_{\mathbf{R}},\#I_{\mathbf{R}},\#N_{\mathbf{R}})$ collecting the three counting functions. The margin function is defined as before:
@@
    If the domain of $F$ is $\mathsf{WOSN}$, then $F$ is margin-based if and only if $F$ satisfies Tiebreaking Compensation, Comparable Compensation, and Neutral Indifference.
@@
If a profile contains a ranking with nonempty side-noncomparability class $S$, duplicate the profile, so we have two copies of the ranking, and then linearize $S$ in opposite ways above and below the other candidates. By Homogeneity and Comparable Compensation, this does not change the output of $F$. In this way, we can change any two initial profiles into profiles with no side-noncomparability. Then for part 1 we reason exactly as in the proof of Theorem 2.9 using Preferential Equality, and for part 2 we reason exactly as in the proof of Theorem 3.8.2 using Tiebreaking Compensation and Neutral Indifference.∎